# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v2

@@ III.2 Reversible asymptotic conversion: Proof of the Marvian-Spekkens conjecture
$\left(C_{\mathfrak{g}}(\psi)\right)_{kl}$ $\coloneqq\frac{1}{2}\braket{\psi}{(X_{k}X_{l}+X_{l}X_{k})}{\psi}$
[…]
@@
[…]
@@ III.3 Irreversibility in asymptotic conversion
Condition (B) is easily violated when $\dim G>1$. As a simple illustration, let us analyze $G=SU(2)$ and its unitary representation $e^{\mathrm{i}\sum_{i=x,y,z}\theta^{i}J_{i}}$ with spin operators $J_{i}$. We adopt the $z$-axis as a quantization axis and denote simultaneous eigenstates of $J^{2}\coloneqq J_{x}^{2}+J_{y}^{2}+J_{z}^{2}$ and $J_{z}$ by |$\ket{j,m}$. To simplify the argument, we consider pure states |$\ket{\psi}$ and |$\ket{\phi}$ in a subspace spanned by the highest-weight states {|$\{\ket{j,j}\}_{j=0,\frac{1}{2},1,\frac{3}{2},\cdots,K}$, where $K$ is an integer or a half-integer introduced to make the Hilbert space finite-dimensional. Following the notation in [2], we define
$\mathcal{J}$ $\coloneqq\textstyle\sum_{j=0,\frac{1}{2},1,\frac{3}{2},\cdots,K}j\ket{j,j}\bra{j,j},$ (13)
$\mathcal{M}(\psi)$ $\coloneqq 2\braket{\psi}{\mathcal{J}}{\psi},$ (14)
@@
and similarly for $\phi$, condition (B) holds only of $\mathcal{M}(\psi)/\mathcal{M}(\phi)=\mathcal{V}(\psi)/\mathcal{V}(\phi)$. This can also be explicitly confirmed from the conversion rate $R(\psi\to\phi)=\min\left\{\frac{\mathcal{M}(\psi)}{\mathcal{M}(\phi)},\frac{\mathcal{V}(\psi)}{\mathcal{V}(\phi)}\right\}$ calculated from Eq. (6) for pure states satisfying $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$. We remark that this conversion rate is consistent with the prior result on $SU(2)$ in Theorem 24 in [2], where the conversion rate is studied for a restricted set of pure states.
@@ III.4 Finite groups
Here, we analyze the approximate asymptotic conversion theory in RTA for a finite group. As proven in [34], the conversion rate diverges for a finite group. We show that this fact can also be derived from our formula in Eq. (5) by lifting a finite group to a compact Lie group by appending a trivial phase, though a finite group itself is not a Lie group.
@@ IV.1 Monotonicity of QGT in one-shot conversion
$m_{f}^{-1}(\mathcal{R}_{\rho},\mathcal{L}_{\rho})(B)\coloneqq\sum_{k,l;m_{f}(p_{k},p_{l})>0}\frac{\braket{l}{B}{k}}{m_{f}(p_{k},p_{l})}\ket{l}\bra{k},$ (19)
where we used the eigenvalue decomposition of a state $\rho=\sum_{k}p_{k}\ket{k}\bra{k}$ |. For a linear operator $A$, we denote $\|A\|_{f,\rho}\coloneqq\sqrt{\braket{A,A}_{f,\rho}}$.
@@
$V(\psi,O)\coloneqq\braket{\psi}{O(I-\psi)O^{\dagger}}{\psi}$ (22)
for a pure state $\psi$ and a linear operator $O$. Here, following a standard convention, we described a pure state and a unit vector by the same Greek letter, i.e., $\psi=\ket{\psi}\bra{\psi}$ |. Note that when $O$ is Hermitian, $V$ is equivalent to the ordinary variance.
@@
In literature, the proof of the monotonicity of monotone metric is sometimes provided under the assumption that $\rho$ is invertible. Lemma 3 is a special case of Theorem 6.1 in [68], which is valid not only for invertible states but also non-invertible states. The details of the proof of Lemma 3 and its relation to Theorem 6.1 in [68] are presented in Appendix B.1.
@@ IV.3 Monotonicity of QGT in asymptotic conversion
Let $U,U^{\prime}$ be projective unitary representations of a Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. For any pure states |$\ket{\psi}\in\mathcal{H}$ and |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$, it holds
@@ V QGT determines asymptotic conversion rate
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. For any pure states |$\ket{\psi}\in\mathcal{H}$ and |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$ such that $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$, it holds
@@
In the proof of Proposition 6, the results in QLAN on pure-state unitary models [56, 57] play a crucial role. This can be intuitively understood through the following lemma, which provides an alternative yet equivalent characterization of the convertibility in RTA using quantum channels that are not assumed to be $G$-covariant:
@@
Lemma 7 implies that $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ if and only if such that
@@ V.1 Quantum local asymptotic normality
For a more detailed explanation, let us introduce several notations in [56]. For a pure state |$\ket{\psi}$ on a qudit system, we define a matrix $C:\mathbb{C}^{m}\to\mathbb{C}^{d-1}$ by $(C)_{ki}\coloneqq-\mathrm{i}\braket{k}{X_{i}}{\psi}$, where {|$\{\ket{k}\}_{k=1}^{d-1}$ denotes an orthonormal basis for the subspace orthogonal to |$\ket{\psi}$. For $\bm{z}=(z_{1},\cdots,z_{d-1})^{\top}\in\mathbb{C}^{d-1}$, we denote a $(d-1)$-mode coherent state with coherence amplitude $\bm{z}$ by |$\ket{\bm{z}}\coloneqq\left(\bigotimes_{i=1}^{d-1}e^{z_{i}a_{i}^{\dagger}-z_{i}^{*}a_{i}}\ket{0}_{i}\right)$, where $a_{i}^{\dagger},a_{i}$ and |$\ket{0}_{i}$ denote the creation and annihilation operator and the vacuum state for the $i$th mode. It is proven [57] that there exist sequences of quantum channels $\{\mathcal{T}_{N}\}_{N}$ and $\{\mathcal{S}_{N}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{T}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi)^{\otimes N}\right),\ket{C\bm{u}}\bra{C\bm{u}}\right)$
$=\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi)^{\otimes N},\mathcal{S}_{N}\left(\ket{C\bm{u}}\bra{C\bm{u}}\right)\right)$
@@
As shown in Eq. (40), $\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi)\right)^{\otimes N}$ is reversibly asymptotically convertible to the Gaussian-shift model |$\ket{C\bm{u}}$, implying that the matrix $C$ characterizes the asymptotic properties of the i.i.d. statistical model $\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi)\right)^{\otimes N}$. Such a matrix $C$ is not uniquely determined since there remains the freedom in the choice of an orthonormal basis {|$\{\ket{k}\}_{k=1}^{d-1}$ for the subspace orthogonal to |$\ket{\psi}$. Adopting another orthonormal basis, the matrix $C$ is transformed to $C\mapsto\tilde{C}\coloneqq VC$, where $V$ denotes a $\mathbb{C}^{d-1}\times\mathbb{C}^{d-1}$ unitary matrix. By using the polar decomposition theorem, for matrices $C$ and $C^{\prime}$ of the same size, we find that $\tilde{C}=VC$ holds for some unitary matrix $V$ if and only if $C^{\dagger}C=\tilde{C}^{\dagger}\tilde{C}$. This fact implies that the essential characteristic of a statistical model is $C^{\dagger}C$, rather than the matrix $C$ itself.
@@
[…]
$=\braket{\psi}{X_{i}(I-\psi)X_{j}}{\psi},$ (41)
@@ V.3 Proof of achievability of the conversion rate
$p^{\mathrm{succ.}}(\delta)\coloneqq$ $\inf_{g\in G}\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes n})\chi_{G_{\mathrm{succ.}}^{(g,\delta)}}(\hat{g}),$ (44)
where $\mu_{G}$ is the Haar measure on the Lie group $G$, and $\chi_{A}$ denotes the indicator function of a set $A$. Note that the above success probability is well-defined since $G_{\mathrm{succ.}}(\delta)$ is a measurable set, as shown in Appendix C.3.
@@
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let |$\ket{\psi}\in\mathcal{H}$ and |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$ be pure states. We define $\psi_{g}\coloneqq\mathcal{U}_{g}(\psi)$ and $\phi_{g}\coloneqq\mathcal{U}^{\prime}_{g}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi_{g}}$ and $\mathcal{Q}^{\phi_{g}}$. If $r>0$ satisfies $\mathcal{Q}^{\psi_{g}}\geq r\mathcal{Q}^{\phi_{g}}$ for all $g\in G$, then, there exists a sequence of quantum channels $\{\mathcal{E}^{(g)}_{N}\}_{N}$ such that the conversion error
@@
×$\quad\times\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$\geq\inf_{g\in G}\int\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})\chi_{G_{\mathrm{succ.}}^{(g,N^{-1/2+\epsilon})}}(\hat{g})$
×$\quad\times\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
@@ VI.2 Sublinear additional resource
For this sequence of the channels, by following the arguments from Eq. (48) to Eq. (52) in the paper, we complete the proof of the direct part, i.e.,
@@ VII.1 Distillation of asymmetry
We start the analysis by relating $\mathcal{S}^{\rho}$ with the monotone metric with an operator monotone function $f_{q}(x)$ given in Eq. (20). By using the eigenvalue decomposition $\rho=\sum_{k}p_{k}\ket{k}\bra{k}$ | of the state $\rho$, we have
$\|\mathrm{i}[O,\rho]\|_{f_{q},\rho}^{2}=\sum_{\underset{(1-q)p_{l}+qp_{k}>0}{k,l;}}\frac{(p_{l}-p_{k})^{2}}{(1-q)p_{l}+qp_{k}}|\braket{l}{O}{k}|^{2}.$ (65)
When $q\to 1^{-}$, the above sum contains divergent terms proportional to $1/(1-q)$ for $(k,l)$ such that $p_{l}>0$ and $p_{k}=0$. Therefore, we find $\lim_{q\to 1^{-}}f_{q}(0)\|\mathrm{i}[O,\rho]\|_{f_{q},\rho}^{2}=\sum_{l;p_{l}>0}\sum_{k;p_{k}=0}p_{l}\braket{l}{O}{k}\braket{k}{O^{\dagger}}{l}$, i.e.,
@@
We now prove the monotonicity of $\mathcal{S}^{\rho}$ under $G$-covariant operation. Suppose a state $\rho$ is convertible to $\sigma$ via a $G$-covariant channel without error. From Eq. (28), we have
@@
Taking the limit of $q\to 1^{-}$ after multiplying the above inequality by $f_{q}(0)=1-q>0$, Eq. (67) implies $\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}\geq\bm{\gamma}^{\dagger}\mathcal{S}^{\sigma}\bm{\gamma}$. Since this inequality holds for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we get
@@
We remark that in a special case where $[\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ holds for some $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we have $\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}=0$. As an immediate corollary of the above theorem, we therefore get the following:
@@
If there exists a vector $\bm{\gamma}\in\mathbb{C}^{\dim G}$ satisfying $[\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ and $\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}\neq 0$, then $\mathcal{A}_{\mathrm{d}}(\rho)=0$.
This corollary implies that for a typical state, which has full rank, $\mathcal{A}_{\mathrm{d}}(\rho)=0$ holds for any asymmetric pure state $\phi$. This result is consistent with a result in prior research [19] on time-translation asymmetry, where the distillable coherence, defined as the optimal conversion rate from a state $\rho$ to a pure state $\phi$ given by a superposition of the ground and excited state of a Hamiltonian $H$, is shown to be zero when $[\Pi_{\rho},H]=0$. We remark that although a state $\rho$ satisfying the condition in Corollary 13 has zero distillable asymmetry, it does not necessarily implies that $\rho$ is useless in asymptotic conversion. For example, while a full-rank state has zero distillable asymmetry, a sublinear number of such states may be used to enhance the conversion rate by removing the restriction imposed by symmetry subgroups in asymptotic conversion among pure states, as stated in Theorem 11.

### v3

@@ III.2 Reversible asymptotic conversion: Proof of the Marvian-Spekkens conjecture
$\left(C_{\mathfrak{g}}(\psi)\right)_{kl}$ $\coloneqq\frac{1}{2}\braket{\psi|(X_{k}X_{l}+X_{l}X_{k})|\psi}$
$\quad-\braket{\psi|X_{k}|\psi}\braket{\psi|X_{l}|\psi}.$ (16)
@@
    $\braket{\psi|L|\psi}=R(\psi\to\phi)\braket{\phi|L|\phi}$ for any element $L$ in the representation of the commutator subalgebra $\mathrm{i}[\mathfrak{g},\mathfrak{g}]$.
@@ III.3 Irreversibility in asymptotic conversion
$\mathcal{J}$ $|,\displaystyle\coloneqq\textstyle\sum_{j=0,\frac{1}{2},1,\frac{3}{2},\cdots,K}j\ket{j,j}\bra{j,j}$, (17)
$\mathcal{M}(\psi)$ $\coloneqq 2\braket{\psi|\mathcal{J}|\psi},$ (18)
$\mathcal{V}(\psi)$ $\coloneqq 4(\braket{\psi|\mathcal{J}^{2}|\psi}-\braket{\psi|\mathcal{J}|\psi}^{2}).$ (19)
@@
and similarly for $\phi$, condition (B) holds only if $\mathcal{M}(\psi)/\mathcal{M}(\phi)=\mathcal{V}(\psi)/\mathcal{V}(\phi)$. This can also be explicitly confirmed from the conversion rate $R(\psi\to\phi)=\min\left\{\frac{\mathcal{M}(\psi)}{\mathcal{M}(\phi)},\frac{\mathcal{V}(\psi)}{\mathcal{V}(\phi)}\right\}$ calculated from Eq. (7) for pure states satisfying $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$. We remark that this conversion rate is consistent with the prior result on $SU(2)$ in Theorem 24 in [12], where the conversion rate is studied for a restricted set of pure states.
@@ III.4 Finite groups
Here, we analyze the approximate asymptotic conversion theory in RTA for a finite group. As proven in [59], the conversion rate diverges for a finite group. We show that this fact can also be derived from our formula in Eq. (6) by lifting a finite group to a compact Lie group by appending a trivial phase, though a finite group itself is not a Lie group  [^1] Formally, a finite group can be regarded as a Lie group since it can be viewed as a zero-dimensional smooth manifold. However, in our terminology, we consider Lie groups to have a dimension greater than zero..
@@ IV.1 Monotonicity of QGT in one-shot conversion
$|,\displaystyle m_{f}^{-1}(\mathcal{R}_{\rho},\mathcal{L}_{\rho})(B)\coloneqq\sum_{k,l;m_{f}(p_{k},p_{l})>0}\frac{\braket{l|B|k}}{m_{f}(p_{k},p_{l})}\ket{l}\bra{k}$, (23)
where we used the eigenvalue decomposition of a state $|\rho=\sum_{k}p_{k}\ket{k}\bra{k}$. For a linear operator $A$, we denote $\|A\|_{f,\rho}\coloneqq\sqrt{\braket{A,A}_{f,\rho}}$.
@@
$V(\psi,O)\coloneqq\braket{\psi|O(I-\psi)O^{\dagger}|\psi}$ (26)
for a pure state $\psi$ and a linear operator $O$. Here, following a standard convention, we described a pure state and a unit vector by the same Greek letter, i.e., $|\psi=\ket{\psi}\bra{\psi}$. Note that when $O$ is Hermitian, $V$ is equivalent to the ordinary variance.
@@
In literature, the proof of the monotonicity of monotone metric is sometimes provided under the assumption that $\rho$ is invertible. Lemma 3 is a special case of Theorem 6.1 in [82], which is valid not only for invertible states but also for non-invertible states. The details of the proof of Lemma 3 and its relation to Theorem 6.1 in [82] are presented in Appendix B.1.
@@ IV.3 Monotonicity of QGT in asymptotic conversion
[…]
@@ V QGT determines asymptotic conversion rate
[…]
@@
In the proof of Proposition 6, the results in quantum local asymptotic normality (QLAN) on pure-state unitary models [86, 87] play a crucial role. This can be intuitively understood through the following lemma, which provides an alternative yet equivalent characterization of the convertibility in RTA using quantum channels that are not assumed to be $G$-covariant:
@@
Lemma 7 implies that $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ if and only if there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that
@@ V.1 Quantum local asymptotic normality
[…]
@@
[…]
@@
$\left(C^{\dagger}C\right)_{ij}$ $=\sum_{k=1}^{d-1}\braket{\psi|X_{i}|k}\braket{k|X_{j}|\psi}$
$=\braket{\psi|X_{i}(I-\psi)X_{j}|\psi},$ (45)
@@ V.3 Proof of achievability of the conversion rate
[…]
where $\mu_{G}$ is the Haar measure on the Lie group $G$, and $\chi_{A}$ denotes the indicator function of a set $A$. Note that the above success probability is well-defined since $G_{\mathrm{succ.}}^{(g,\delta)}$ is a measurable set, as shown in Appendix C.3.
@@
[…]
@@
$\quad\times\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
[…]
$\quad\times\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
@@ VI.2 Sublinear additional resource
For this sequence of the channels, by following the arguments from Eq. (52) to Eq. (56), we complete the proof of the direct part, i.e.,
@@ VII.1 Distillation of asymmetry
We start the analysis by relating $\mathcal{S}^{\rho}$ with the monotone metric with an operator monotone function $f_{q}(x)$ given in Eq. (24). By using the eigenvalue decomposition $|\rho=\sum_{k}p_{k}\ket{k}\bra{k}$ of the state $\rho$, we have
$\|\mathrm{i}[O,\rho]\|_{f_{q},\rho}^{2}=\sum_{\underset{(1-q)p_{l}+qp_{k}>0}{k,l;}}\frac{(p_{l}-p_{k})^{2}}{(1-q)p_{l}+qp_{k}}|\braket{l|O|k}|^{2}.$ (69)
When $q\to 1^{-}$, the above sum contains divergent terms proportional to 1/(1q)1/(1-q) for $(k,l)$ such that $p_{l}>0$ and $p_{k}=0$. Therefore, we find $\lim_{q\to 1^{-}}f_{q}(0)\|\mathrm{i}[O,\rho]\|_{f_{q},\rho}^{2}=\sum_{l;p_{l}>0}\sum_{k;p_{k}=0}p_{l}\braket{l|O|k}\braket{k|O^{\dagger}|l}$, i.e.,
@@
We now prove the monotonicity of $\mathcal{S}^{\rho}$ under $G$-covariant operation. Suppose a state $\rho$ is convertible to $\sigma$ via a $G$-covariant channel without error. Multiplying Eq. (32) by $f_{q}(0)=1-q>0$ yields
@@
Taking the limit of $q\to 1^{-}$, Eq. (71) implies $\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}\geq\bm{\gamma}^{\dagger}\mathcal{S}^{\sigma}\bm{\gamma}$. Since this inequality holds for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we get
@@
We remark that in a special case where $\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ holds for some $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we have $\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}=0$. As an immediate corollary of the above theorem, we therefore get the following:
@@
If there exists a vector $\bm{\gamma}\in\mathbb{C}^{\dim G}$ satisfying $\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ and $\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}\neq 0$, then $\mathcal{A}_{\mathrm{d}}(\rho)=0$.
This corollary implies that for a typical state with full rank, $\mathcal{A}_{\mathrm{d}}(\rho)=0$ holds for any asymmetric pure state $\phi$. This result is consistent with a result in prior research [23] on time-translation asymmetry, where the distillable coherence, defined as the optimal conversion rate from a state $\rho$ to a pure state $\phi$ given by a superposition of the ground and excited state of a Hamiltonian $H$, is shown to be zero when $\Pi_{\rho},H]=0$. We remark that although a state $\rho$ satisfying the condition in Corollary 13 has zero distillable asymmetry, it does not necessarily implies that $\rho$ is useless in asymptotic conversion. For example, while a full-rank state has zero distillable asymmetry, a sublinear number of such states may be used to enhance the conversion rate by removing the restriction imposed by symmetry subgroups in asymptotic conversion among pure states, as stated in Theorem 11.